# Brain-inspired Large Language Model

### v3

@@ Abstract
This paper reports the first brain-inspired large language model (BriLLM). This is a non-Transformer, non-GPT, non-traditional machine learning input-output controlled generative language model. The model is based on the Signal Fully-connected flowing (SiFu) definition on the directed graph in terms of the neural network, and has the interpretability of all nodes on the graph of the whole model, instead of the traditional machine learning model that only has limited interpretability at the input and output ends. In the language model scenario, the token is defined as a node in the graph. A randomly shaped or user-defined signal flow flows between nodes on the principle of ”least resistance” along paths. The next token or node to be predicted or generated is the target of the signal flow. As a language model, BriLLM theoretically supports infinitely long $n$-gram models when the model size is independent of the input and predicted length of the model. The model’s working signal flow provides the possibility of recall activation and innate multi-modal support similar to the cognitive patterns of the human brain. At present, we released the first BriLLM version in Chinese, with 4000 tokens, 32-dimensional node width, 16-token long sequence prediction ability, and language model prediction performance comparable to GPT-1. More computing power will help us explore the infinite possibilities depicted above.  [^1] Code at: https://github.com/brillm05/BriLLM0.5  [^2] Model at: https://huggingface.co/BriLLM/BriLLM0.5
@@ 1 Introduction
Large language models (LLMs) are igniting the prospect of AGI (artificial general intelligence). However, even SOTA LLMs are still in terms of Transformer architecture and GPT training scheme unlikely to laugh at the final termination of AGI due to the huge difficulties in their scalability and interpretability, let alone the way Transformer or GPT-based LLM works is a far cry from the human brain, the alternative intelligence machine already existing in the nature for millions of years, showing how a true AGI must be.
The Transformer (Vaswani et al., 2017) has been a fundamental and indispensable framework for building SOTA LLM backbones. Although Transformers have demonstrated remarkable generalization capabilities across diverse tasks and scalability to achieve higher intelligence, the quadratic computational complexity of the attention mechanism over input sequences poses significant efficiency challenges, particularly for long sequences. This computational bottleneck has spurred research into more efficient attention variants, such as linear attention mechanisms, and RNN-like Transformers. While these studies focus on preserving model performance and lowering computational costs, they merely mitigate the issue without resolving the computational bottleneck at its core, since they remain dependent on attention-based mechanisms or attention variants.
Furthermore, the Transformer architecture exhibits limited parameter-level interpretability due to its complex self-attention mechanisms and opaque parameter interactions, a characteristic that renders it functionally analogous to a black-box system. Many studies endeavor to reveal the black box by interpreting the intrinsic mechanism of self-attention or enhancing the interpretability of the model through visualization, attribution methods, and probing tasks. However, the complicated interaction of attention between hidden states still remains poorly understood.
@@ 2 SiFu Mechanism
Inspired by the working mode of the brain, we propose Signal Fully-connected Flowing (SiFu) on the Directed Graph, a novel input-output stream control mechanism for machine learning, serving as the core design of BriLLM. As shown in Figure 1, SiFu model is a graph composed of multiple nodes, which are sparsely activated and utilize tensors to transmit a nominal signal. Each node (ideally, a layer of neurons) represents a certain concept or word, e.g., a noun, a verb, etc. Each edge models the relationship between every node pair. The signal is transmitted by the magnitude of the energy. The energy will be strengthened, i.e., maximized, if it is in the right route. Or, at least, the right path always keeps the maximal energy for the transmitted signal. Each node is sequentially activated in terms of the maximized energy. Route or path is determined in a competitive way, i.e., the next node will be activated only if the energy can be maximally delivered in this node.
@@
Figure 1: An illustration of SiFu Directed Graph.
@@
For example, as shown in Figure 1, the path “dog $\to$ love $\to$ meat” has the largest energy. As shown in Figure 1, the correct sequence should yield the largest energy. For example, to calculate the loss for the sequence “love $\to$ meat”: multiple negative samples in the vocabulary, such as “apple” and “iron,” are selected. Energy tensors are computed for both the ground-truth node (“meat”) and negative nodes (“apple”, “iron”). The cross-entropy loss maximizes the energy associated with the node “meat” while minimizing energies from the negative nodes.
@@ 3 BriLLM Formulation
BriLLM implements SiFu neural network for language modeling, as shown in Figure 3. Each token in the vocabulary is modeled as a node, which is defined by a hidden layer of neurons in the neural network. Similarly to other sequence transduction models, we convert each token into a learned embedding vector $x_{i}\in\mathbb{R}^{d_{model}}$. An edge connecting token $u$ and $v$ is modeled as a fully-connected matrix $W_{u,v}\in\mathbb{R}^{d_{model}\times d_{model}}$ and a bias $b_{u,v}\in\mathbb{R}^{d_{model}}$. Two fully-connected matrices $W_{u,v}$ and $W_{v,u}$ play the roles of the bidirectional edges between tokens. The signal tensors are fitted into matrices. The forward process begins with an initial signal shape:
$e_{1}=\mathbf{1}_{n}=[1,1,\dots,1]^{\top}\in\mathbb{R}^{d_{model}}$ (1)
Then an expanded signal tensor $\mathcal{E}_{i}\in\mathbb{R}^{d_{model}}$ is computed as a linear weighted sum of previous signals using learnable weights $w\in\mathbb{R}^{maxlen}$:
$\mathcal{W}$ $=\mathop{{\rm softmax}}(w_{1:i})$ (2)
$\mathcal{E}_{i}$ $=\sum_{k=1}^{i}\mathcal{W}_{k}e_{k},$ (3)
where $\mathcal{W}$ represents the softmax-normalized weights.
The signal flowing from token $x_{i}$ to all other tokens $e_{i+1,v}\in\mathbb{R}^{d_{model}}$ will be computed:
$e_{i+1,v}={\rm GeLU}(W_{x_{i},v}\mathcal{E}_{i}+b_{x_{i},v}+PE_{i}),$
where $PE$ represents the sine and cosine positional encoding.
During decoding, the next token $x_{i+1}$ is determined by the edge with the largest energy:
$x_{i+1}$ $=\mathop{\arg\max}_{v}\|{e_{i+1,v}}\|_{2}$ (4)
$e_{i+1}$ $=\mathop{\max}\|{e_{i+1,v}}\|_{2}$
The L2 norm of the signal tensor represents its energy magnitude.
@@
To train a sample in BriLLM, every time we build an individual common neural network to perform the regular BP training. This network consists of two parts, in which the front part connects all input nodes (i.e., tokens), then it follows the rear parts which connect all possible paths in order. At last, a softmax layer collects all paths’ energy tensors to indicate the right path with a 0-1 ground truth vector. We adopt a cross-entropy loss for training.
@@ 4 Experiments
We released BriLLM-Chinese and BriLLM-English model.
@@
Dataset
For BriLLM-Chinese and BriLLM-English, we use the Chinese and English versions of Wikipedia respectively, each containing over 100M tokens. We truncate the long sentences into small sentences with a maximum length of 32. We select a vocabulary of 4,000 tokens consisting of the most frequently used Chinese characters.
@@
BriLLM is a sparse model, where the occurrence probability of most bigrams is very low or even zero, allowing us to leverage this characteristic for sparse training. We set the connection weights corresponding to low-frequency bigrams (those not appearing in the training set) to be shared and update them randomly. This approach reduces the model size to approximately 10% of the original while significantly accelerating the training speed.
Implementation Details.
BriLLM is implemented using PyTorch. It uses sine and cosine positional encoding, GeLU as the activation function, cross-entropy loss for next-token prediction, and an embedding size of $d_{model}=32$. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The original model size is about $512+4000*4000*(32*32+32)\approx 16B$, but after applying sparse training, the actual size is reduced to around $2B$. We trained our models on one machine with 8 NVIDIA A800 GPUs for 1.5k steps. The training loss is shown in Figure 4.
@@
$n$ is the sequence length, $v$ is the vocabulary size, and $d$ is the representation dimension. The computational complexity is $O(n\cdot v\cdot d^{2})$.
@@
Table 1 presents some of the decoding results, including both training samples and test samples.
@@
Table 1: Case study of decoding results.
@@ 5 Conclusion, Limitation and the Future
Currently, due to our quite limited computational power for this work, we just reach an early model checkpoint with a moderate hyperparameter setting on the Chinese language. However, the current released model has demonstrated promising performance compared to GPT-1 (Radford et al., 2018). Meanwhile, so limited computational resource puts huge obstacles to let us explore much more potential over BriLLM.
Our existing BriLLM implementation has a size of $d_{model}\times d_{model}\times n\times n$, where $n$ is the number of tokens (nodes) and $d_{model}$ is embedding size. This quadratically increasing model size is indeed inconvenience. However, as most model parameters come from the fully-connected matrices, it is possible to adopt a sort of sparse representation or shared parameters for those less active tokens, i.e., set a default non-updated matrix for all these inactive tokens. Our preliminary estimation show such strategy may save up to 90% parameters for BriLLM.
@@
The last but not the least issue we need to explore about BriLLM is the possibility of supervised finetuning (SFT) like LLM. Note that as BriLLM does not need to resize the model for any sized input or output sequences and the size BriLLM has to be quadratically correlated to the embedding size and token numbers, it is not in an advantageous position when the model sizes are the same ’small’ or moderate as LLM. As we reported in this paper, a 16B BriLLM (our current released checkpoint) only gives comparable performance as 0.1B GPT-1. Thus, we have reasons to speculate that BriLLM has a very high emergent ability threshold. What’s more, now we even do not know how to do SFT over BriLLM, which leaves a big future work.

### v4

# BriLLM: Brain-inspired Large Language Model

@@ Abstract
This paper reports the brain-inspired large language model (BriLLM). This is a non-Transformer, non-GPT, non-traditional machine learning input-output controlled generative language model. The model is based on the Signal Fully-connected flowing (SiFu) definition on the directed graph in terms of the neural network, and has the interpretability of all nodes on the graph of the whole model, instead of the traditional machine learning model that only has limited interpretability at the input and output ends. In the language model scenario, the token is defined as a node in the graph. A randomly shaped or user-defined signal flow flows between nodes on the principle of "least resistance" along paths. The next token or node to be predicted or generated is the target of the signal flow. As a language model, BriLLM theoretically supports infinitely long $n$-gram models when the model size is independent of the input and predicted length of the model. The model’s working signal flow provides the possibility of recall activation and innate multi-modal support similar to the cognitive patterns of the human brain. At present, we released the first BriLLM versions in Chinese and English, with 4000 tokens, 32-dimensional node size, 32-token sequence prediction ability, model sizes around 2B and 1B respectively, bringing language model prediction performance comparable to GPT-$1$Code at: https://github.com/brillm05/BriLLM0.5  [^2] Model at: https://huggingface.co/BriLLM/BriLLM0.5.
@@ 1 Introduction
Large language models (LLMs) are igniting the prospect of AGI (artificial general intelligence). However, even SOTA LLMs are still in terms of Transformer architecture and GPT training scheme unlikely to laugh at the final termination of AGI due to the huge difficulties in their scalability and interpretability, let alone the way Transformer or GPT-based LLM works is a far cry from the human brain, the alternative intelligence machine already existing in nature for millions of years, showing how a true AGI must be.
The Transformer (Vaswani et al., 2017) has been a fundamental and indispensable framework for building SOTA LLM backbones. Although Transformers have demonstrated remarkable generalization capabilities across diverse tasks and scalability to achieve higher intelligence, the quadratic computational complexity of the attention mechanism over input sequences poses significant efficiency challenges, particularly for long sequences. This computational bottleneck has spurred research into more efficient attention variants, such as linear attention mechanisms, and RNN-like Transformers. Although these studies focus on preserving model performance and lowering computational costs, they merely mitigate the issue without resolving the computational bottleneck at its core, since they remain dependent on attention-based mechanisms or attention variants.
Furthermore, the Transformer architecture exhibits limited parameter-level interpretability due to its complex self-attention mechanisms and opaque parameter interactions, a characteristic that renders it functionally analogous to a black-box system. Many studies attempt to reveal the black box by interpreting the intrinsic mechanism of self-attention or enhancing the interpretability of the model through visualization, attribution methods, and probing tasks. However, the complicated interaction of attention between hidden states remains poorly understood.
@@ 2 SiFu Mechanism
Inspired by the working mode of the brain, we propose Signal Fully-connected Flowing (SiFu) on the Directed Graph, a novel input-output stream control mechanism for machine learning, serving as the core design of BriLLM. As shown in Figure 1, SiFu model is a graph composed of multiple nodes, which are sparsely activated and utilize tensors to transmit a nominal signal. Each node (ideally, a layer of neurons) represents a certain concept or word, e.g., a noun, a verb, etc. Each edge models the relationship between every pair of nodes. The signal is transmitted by the magnitude of the energy. The energy will be strengthened, i.e., maximized, if it is in the right route. Or, at least, the right path always keeps the maximal energy for the transmitted signal. Each node is sequentially activated in terms of the maximized energy. The route or path is determined in a competitive way, i.e., the next node will be activated only if the energy can be maximally delivered in this node.
@@
Figure 1: An illustration of SiFu Directed Graph (Numbers by the node denote energy scores).
@@
For example, as shown in Figure 1, the path “dog $\to$ love $\to$ meat" has the highest energy. As shown in Figure 1, the correct sequence should yield the highest energy. For example, to calculate the loss for the sequence “love $\to$ meat": multiple negative samples in the vocabulary, such as “apple" and “iron," are selected. Energy tensors are computed for both the ground-truth node (“meat") and negative nodes (“apple", “iron"). A chosen loss function maximizes the energy associated with the node “meat" while minimizing energies from the negative nodes.
@@ 3 BriLLM Formulation
BriLLM implements SiFu neural network for language modeling, as shown in Figure 3. Each token in the vocabulary is modeled as a node, which is defined by a hidden layer of neurons with GeLU activation function and a bias $b\in\mathbb{R}^{d_{node}}$, where $d_{node}$ denotes node size, i.e., how many neuron in a node. An edge connecting nodes $u$ and $v$ is modeled as a fully-connected matrix $W_{u,v}\in\mathbb{R}^{d_{node}\times d_{node}}$. Two fully-connected matrices $W_{u,v}$ and $W_{v,u}$ play the roles of the bidirectional edges between nodes. The signal tensors are fitted into matrices. The forward process begins with an initial signal shape:
$e_{0}=[1,1,\dots,1]^{\top}\in\mathbb{R}^{d_{node}}$ (1)
Suppose we have a token sequence, $u_{1},...,u_{L-1},v_{predict}$, as a training sample. When the signal flows from a node $u_{i}$ to its next node $u_{i+1}$, the energy tensor $e_{{i+1}}\in\mathbb{R}^{d_{node}}$ will be computed:
$e_{{i+1}}=\left\{\begin{aligned} {\rm GeLU}(W_{u_{i},u_{i+1}}e_{i}+b_{u_{i},u_{i+1}}+PE_{i})\;\;{\rm if}\;\;i>0&\\ {\rm GeLU}(e_{0}+b_{u_{1}}+PE_{0})\;\;{\rm if}\;\;i=0\end{aligned}\right.$
where $PE$ represents the sine and cosine positional encoding. Note that we have an edge sensitive bias setting for each node taking inputs. When a node starts a sequence, there is no edge difference, i.e., node $u_{1}$ has an edge independent bias $b_{u_{1}}$ in this case.
To predict a token (node), an expanded signal tensor $\mathcal{E}_{i}\in\mathbb{R}^{d_{node}}$ is computed as a linear weighted sum of previous signals using learnable weights $w\in\mathbb{R}^{L-1}$:
$\mathcal{W}$ $=\mathop{{\rm softmax}}(w_{1:L-1})$ (2)
$\mathcal{E}_{L-1}$ $=\sum_{k=1}^{L-1}\mathcal{W}_{k}e_{k},$ (3)
where $L$ is sequence length and $\mathcal{W}$ represents the softmax-normalized weights. The learnable weights $w$ let the predicted token pay “attention" to all previous tokens other than the directly connected one.
At last, the final energy tensor for next token prediction is computed by:
$E_{u,v}={\rm GeLU}(W_{u_{L-1},v}\mathcal{E}_{L-1}+b_{u_{L-1},v}+PE_{L-1}),$
During inference, the model finds the right predicted token $v_{predict}$ which has the largest energy:
$v_{predict}$ $=\mathop{\arg\max}_{v}\|{E_{u,v}}\|_{2}$ (4)
where the L2 norm of the signal tensor computes its energy score or magnitude.
@@
To train a token sequence sample in BriLLM, every time we build an individual common neural network to perform the regular BP training. This network consists of two parts, in which the front part connects all input nodes (i.e., tokens), then it follows the rear parts which connect all possible paths in order. At last, a softmax layer collects all paths’ energy tensors to indicate the right path with a 0-1 ground truth vector. We adopt a cross-entropy loss for training.
@@ 4 Experiments
We released BriLLM-Chinese and BriLLM-English models.
@@
Datasets
For BriLLM-Chinese and BriLLM-English, we use the Chinese and English versions of Wikipedia respectively, each containing over 100M tokens. We truncate the long sentences into small sentences with a maximum length of 32. We select a vocabulary of 4,000 tokens for both languages.
Implementation Details.
BriLLM is implemented using PyTorch. It uses sine and cosine positional encoding, GeLU as the activation function, cross-entropy loss for next-token prediction, and a node size of $d_{node}=32$. We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The original model size is about $512+4000*4000*(32*32+32)\approx 16B$. We trained our models on one machine with 8 NVIDIA A800 GPUs for 1.5k steps. The training loss is shown in Figure 4.
@@
BriLLM enables sparse training, where the occurrence probability of most bigrams is very low or even zero, allowing us to leverage this characteristic for sparse training. We set the connection weights corresponding to low-frequency bigrams (those not appearing in the training set) to be shared and update them randomly. After applying sparse training, the actual size of BriLLM-Chinese and BriLLM-English is reduced to 2B and 1B, respectively, as shown in Table 1. This approach reduces the model size to approximately 10% of the original while significantly accelerating the training speed.
Table 1: Model sizes before and after sparse training.
BriLLM-Chinese BriLLM-English
original 16.90B 16.90B
sparse 2.19B 0.96B
ratio 13.0% 5.7%
@@
Let $L$ be the sequence length, $n$ the vocabulary size, and $d_{node}$ the node size (dimension), then the forward computational complexity of BriLLM is $O(L\cdot n\cdot d_{node}^{2})$.
@@
Tables  2 and  3 present some of the decoding results, including both training samples and test samples for Chinese and English, respectively.
@@
Table 2: Case study of BriLLM-Chinese decoding results.
Input Completion
Training samples
In frogs, the hind legs are larger In frogs, the hind legs are larger than taxation arrangements and terms, misconstd Paris Academy members of Portals
The requirement for the Sun angle was The requirement for the Sun angle was arguments from Intr proposed: documentary directed by employing hundreds reduced by employe 11 September 1972
The English biologist Thomas Henry Huxley The English biologist Thomas Henry Huxley coined World C that ADE XaZul 30 Ars lead singular shipb more smaller im
Physicist Richard Feynman was noted for facility Physicist Richard Feynman was noted for facility in him increasingly holding six countries, misconstd atomic freedom before
Elements heavier than iron were Elements heavier than iron were retreatywriter 10th worked (ital magnitude, misconstd atomic Music freedom
Typically, when an algorithm is associated with Typically, when an algorithm is associated with Achill declaraus, misconceptions presented at Irraditional emotunday Prich
Plants are used as herbs Plants are used as herbs and Earth Day of Portals working on recent years of Portals working on recent genocots only marked serious risk that
The term vestibular The term vestibular at Texas variable Spec struggathological ideal remains the division of value of value cannot be supern2
Knight’s criticism greatly damaged van Knight’s criticism greatly damaged vanand soon to: examples are ’to looked identity said to: accounts reduced by employe
Atlas-Imperial, an American Atlas-Imperial, an American Advideo game), December with Achill declar between 2003, misconstd atomic freedom in
Test samples
The islands have The islands have been cultivated less than form of value and 1969 via the division of value, miscons lead to non-ane rock
The blue whale (Balaenoptera musculus) The blue whale (Balaenoptera musculus) order in him responsibility of Portals working on recent gene 11 September 197
The Vincent Price film, House of Wax The Vincent Price film, House of Waxi theorem approached the sequel strikend across the sequel strikend across
The Jewish Encyclopedia reports, In February The Jewish Encyclopedia reports, In February 11th worked in him increasingly holds reduced by employe 11 September 1972
The Bermuda Triangle The Bermuda Triangle, Azerbaijani official letters) markeditors), highest number of Portals working on recent years, misconcept of
Table 3: Case study of BriLLM-English decoding results.
@@ 5 Conclusion, Limitation and the Future
Currently, due to our quite limited computational power for this work, we just reach early model checkpoints with a moderate hyperparameter setting. However, the current released models have demonstrated promising performance compared to GPT-1 (Radford et al., 2018).
To precisely understand the SiFu learning mechanism or BriLLM, one must realize that their biggest difference from traditional machine learning is that the former supports multiple concurrent multiple input and multiple output streams, while the latter can only physically accept one input at a time while managing one output. We envision an embodied intelligent implementation of BriLLM, where nominal signals can be multiple, and multiple signal streams can propagate independently along different paths without interference inside the BriLLM, guided by the principle of energy maximization, thereby achieving synchronous multiple inputs and outputs. According to the definition of SiFu learning, this means that each signal stream path represents a pathway of thought. This is very similar to how the human brain synchronously engages in multiple thoughts and multitasks, which is difficultly implemented in terms of traditional single input single output stream of machine learning methods.
In addition to the differences and connections between SiFu learning and traditional machine learning, another interesting topic that naturally arises is whether SiFu learning can be considered a form of representation learning in the context of deep learning. However, it seems difficult to draw such a conclusion. Currently, in the implementation of the BriLLM model, the only learnable weight at the most critical node definition is the bias vector $b$. However, $b$ itself does not carry any motivation for representation learning, because according to the original design of SiFu learning, the role of $b$ is merely to filter the same signal flow into different shapes. Therefore, even if we view the bias vector $b$ as some form of embedded representation for a node like deep learning, it is still a very weak form of representation, far from the strong representation forms that are directly and clearly defined in representation learning.
Our current BriLLM implementation has a size of $(n\times(1+d_{node}))^{2}$+$n^{2}\times d_{node}$, where $n$ is the number of tokens (nodes) and $d_{node}$ is the node size. This quadratically increasing model size is indeed inconvenience. However, as most model parameters come from the fully-connected matrices, we have shown that it is possible to adopt a sort of sparse representation or shared parameters for those less active tokens, i.e., set a default non-updated matrix for all these inactive tokens. Our empirical results in Table 1 show such strategy may save up to about 90% or more parameters for BriLLM.
@@
The last but not the least issue we need to explore about BriLLM is the possibility of supervised finetuning (SFT) like LLM. Note that as BriLLM does not need to resize the model for any sized input or output sequences and the size BriLLM has to be quadratically correlated to the node size and token numbers, it is not in an advantageous position when the model sizes are the same ’small’ or moderate as LLM. As we reported in this paper, a 1-2B BriLLM (our current released checkpoints) only gives comparable performance as 0.1B GPT-1. Thus, we have reasons to speculate that BriLLM has a very high emergent ability threshold. What’s more, now we even do not know how to do SFT over BriLLM, which leaves a big future work.